# Did I Faithfully Say What I Thought? Bridging the Gap Between LLMs Neural Activity and Self-Explanations

### v3

@@ 7 Conclusion
This work introduces NeuroFaith, a flexible framework that measures self-NLE faithfulness by comparing explanations against mechanistic analysis of model internal activity. NeuroFaith aligns more closely than existing approaches with the common acceptance of faithfulness. Our key contributions suggest that faithfulness exhibits linear structure in LLM representation space, enabling both fast and reliable detection and enhancement through steering interventions. These findings establish concrete pathways toward more transparent AI systems, opening up new avenues of research on faithfulness. NeuroFaith relies on concept extraction methods that may introduce systematic biases, and circuit selection requires domain expertise or expensive circuit discovery that may limit generalizability. Our analysis focuses on predict-then-explain scenarios; extending to explain-then-predict settings could reveal how explanation faithfulness relates to performance gains (Bhan et al., 2024). Additionally, applying NeuroFaith to chain-of-thought reasoning could provide valuable comparisons with existing CoT faithfulness studies (Lanham et al., 2023; Turpin et al., 2023), advancing our understanding of reasoning faithfulness.
Appendix A Appendix
@@ A Scientific Libraries
We used several open-source libraries in this work: pytorch (Paszke et al., 2019), HuggingFace transformers (Wolf et al., 2020) and sklearn Pedregosa et al. (2011).
@@ C Datasets
Here we provide detailed information about the analyzed datasets. For 2-hop reasoning, we run NeuroFaith on the Wikidata-2-hop dataset (Biran et al., 2024). We first compute task accuracy on the whole dataset, and then sample 1500 accurate and inaccurate prediction for the following of our study. To sample these 1500 instances, we filter 2-hop reasoning questions where relations are subjective (e.g. ”the most notable work of”) or equivocal, potentially leading to numerous possible answers (e.g. ”the work that features” ). We also sample by setting a maximum number of occurrences for generated answers (15), to foster diversity in the input questions. This filter enables to avoid having to many questions where the answer is a country (e.g. USA). We end up with a dataset made of 3000 samples with 1500 accurate and 1500 inaccurate predictions.
For classification, we run NeuroFaith on AGNews (Gulli, 2005), a newspaper article classification and Ledgar (Tuggener et al., 2020), a more challenging critical domain legal document classification dataset. We retrieve the enriched versions from Bhan et al. (2025) with labeled concepts for CAV computation. For AGNews, the classes to predict are ’world’, ”sport”, ”business” and ”science & technology”. For Ledgar, the classes to predict are ”Amendments”, ”Survival”, ”Terminations” and ”Terms”. Each dataset is made of 4000 samples.
@@ D.1 Concept Extraction
Here we provide the prompts used to give instructions to Qwen-3-32b to extract relevant concepts (NeuroFaith step 1). For 2-hop reasoning, concept extraction consists in retrieving the bridge object from the self-NLE. Since the input text as the following structures : $x=(o_{1},r_{1},{\blacktriangle{}},r_{2},\bullet)$, we directly prompt the model to resolve $(o_{1},r_{1},{\blacktriangle{}}$) by grounding its response on the self-NLE only. We structure our prompt following an in-context learning template with two examples differing from the dataset of interest:
  • <<user>>
@@
  • <<Assistant>>
  • Emmanuel Macron
  • <<user>>
  • preprompt + preprompt_example_2
  • <<Assistant>>
  • Ingmar Bergman
  • preprompt + $(o_{1},r_{1},{\blacktriangle{}}$)
  • <<user>>
@@
For classification, we assess having access to a set of relevant concepts $\mathcal{C}$ and labels related to the task of interest. Given a certain concept $c$, we only prompt the model to assess if the concept is present in the self-NLE if the concept was initially present in the input text, making the concept extraction process computationally less expansive. We structure our prompt following an in-context learning template with three examples differing from the dataset of interest:
  • <<user>>
@@
  • <<Assistant>>
  • YES
  • <<user>>
  • preprompt + preprompt_example_2
  • <<Assistant>>
  • YES
  • <<user>>
  • preprompt + preprompt_example_3
  • <<Assistant>>
  • NO
  • <<user>>
@@ D.2 Mechanistic Interpretation
To generate an mechanistic interpretation of an LLM hidden state, the first step is to define the specific neural architectural component and granularity $\mathcal{G}$. Given an index $k$ and a layer $\ell$, a granularity $\mathcal{G}$ refers to a specific point within the model’s computational subgraph where neural activations can be analyzed and from which a local neural interpretation $i^{\ell}_{k}(x)$ is going to be generated. Transformer-based generative LLM can be viewed as a stack of decoder computational blocks (Radford et al., 2018). The information flow through a single layer and from a layer to another can be described with the following equations:
@@
For the 2-hop reasoning instantiation, the interpreter used to decode $f$ hidden states $\left\{h_{k}^{\ell}\right\}$ is Patchscopes Ghandeharioun et al. (2024). We use the prompt ”What is the following? Answer briefly [X,X]” to generate the interpretation where X is a token placeholder to be replaced in the latent space by the hidden state to be interpreted. We replace the placeholder tokens at layers 3 and 4 to get two interpretations per hidden state to decode. The layers to be interpreted by Patchscopes vary depending on the assessed model. We set the index token $k$ as the last one related to $o_{1}$ and focus on the late early layers as in Biran et al. (2024). This way, $\ell\in\left\{{5,6,7,8,9,10,11}\right\}$ for gemma-2-2B, $\ell\in\left\{{8,9,10,11,12,13,14}\right\}$ for gemma-2-9B and $\ell\in\left\{{11,12,13,14,15,16,17}\right\}$ for gemma-2-27B.
@@
[…]
@@
Faithfulness is finally calculated based on concept importance as follows: $F(x,e)=\frac{1}{p}\sum_{i=1}^{p}\mathbf{1}_{\texttt{I}_{\Gamma}(c_{i})>0}$. Below are two distributions of faithfulness for classification of self-NLE generated for AGnews and Ledgar from gemma-2-27b.
In Figure 5, 6, 7, 8, 9, and 10 we plot the concepts sorted by frequency for faithful and unfaithful self-NLE for AGNews and Ledgar for gemma-2-2b and gemma-2-9b for several class predictions. These figures enable to highlight concepts related to either faithful or unfaithful self-NLE. For example, Figure 8 highlights the concept of ”companies” to the class world, which seems to be counterintuitive.
@@
We also plot the density of faithfulness for AGNews and Ledgar for gemma-2-27b in Figure 11 and 12.
@@
Here we provide additional information about layer-wise linear probe performance and similarity between faithfulness vectors and with other AI safety linear vectors. Figures13, 14, 15, 16, 17, 18, 19, 20, 21 show the layer-wise performance of the faithfulness linear probes for 2-hop reasoning, AGNews and Ledgar. As shown in Figure 30, 23 and 31, cosine similarity between task-specific linear faithfulness and AI safety behaviors vectors becomes more pronounced with the size of the model.
@@ E Detailed Taxonomy of Self-NLE in Two-hop Reasoning
  • ($C_{1}$) Complete reasoning failure. If $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\forall(k,\ell)\in\Gamma,o_{2}\notin\tilde{h^{\ell}_{k}}$. Observation: Wrong prediction, incorrect unfaithful explanation, ground-truth bridge object undetected in circuit $\Gamma$. Interpretation: Evidence suggests failure in first-hop reasoning, with neither explanation nor internal representations containing the expected bridge object.
@@
  • ($C_{2}$) Internal-external reasoning mismatch. If $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\exists(k,\ell)\in\Gamma,o_{2}\in\tilde{h^{\ell}_{k}}$. Observation: Wrong prediction, incorrect unfaithful explanation, but ground-truth bridge object detected in circuit $\Gamma$. Interpretation: Model appears to have correct internal knowledge but generates inconsistent explanations, indicating potential reasoning-explanation dissociation because of either deceptiveness or hallucination.
@@
  • ($C_{3}$) Explanation-prediction association. If $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=0$ and $\widehat{o_{2}}=o_{2}$. Observation: Wrong prediction with unfaithful but correct explanation. Interpretation: Model associates correct bridge object with incorrect prediction during explanation generation, suggesting superficial pattern matching without genuinely resolving the first hop of the 2-hop reasoning.
@@
  • ($C_{4}$) First-hop reasoning failure. If $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}\neq o_{2}$. Observation: Wrong prediction with faithful but incorrect explanation. Interpretation: Model consistently follows incorrect reasoning pathway, indicating systematic error in first-hop reasoning or concept misunderstanding.
@@
  • ($C_{5}$) Second-hop reasoning failure. If $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}=o_{2}$. Observation: Wrong prediction with faithful and correct explanation. Interpretation: Model correctly identifies bridge object but fails in second reasoning step, suggesting error occurs after successful first-hop completion.
@@
  • ($C_{6}$) Shortcut learning. If $\widehat{o_{3}}=o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\forall(k,\ell)\in\Gamma,o_{2}\notin\tilde{h^{\ell}_{k}}$. Observation: Correct prediction with unfaithful incorrect explanation, ground-truth bridge object undetected. Interpretation: Evidence suggests direct $o_{1}\rightarrow o_{3}$ association, consistent with shortcut learning behavior that bypasses intermediate reasoning steps.
@@
  • ($C_{7}$) Deceptiveness or hallucination. If $\widehat{o_{3}}=o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\exists(k,\ell)\in\Gamma,o_{2}\in\tilde{h^{\ell}_{k}}$. Observation: Correct prediction with unfaithful incorrect explanation, but ground-truth bridge object detected internally. Interpretation: Model possesses correct internal knowledge but generates deceptive (or hallucinated) explanations, suggesting reasoning-explanation dissociation or alternative reasoning pathways. This case is expected to be rare, otherwise highlighting a case where $f$ is not honest in its self-NLE while ”knowing” the ground truth bridge object, raising a problem in $f$ alignment.
@@
  • ($C_{8}$) Explainer parrot. If $\widehat{o_{3}}=o_{3}$, $F(x,e)=0$ and $\widehat{o_{2}}=o_{2}$. Observation: Correct prediction with unfaithful but correct explanation. Interpretation: Model generates expected explanations without corresponding detectable internal reasoning, suggesting post-hoc explanation generation (i.e. ”explainer parrot” behavior).
@@
  • ($C_{9}$) Alternative reasoning pathway. If $\widehat{o_{3}}=o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}\neq o_{2}$. Observation: Correct prediction with faithful but incorrect explanation. Interpretation: Model uses consistent but non-canonical reasoning pathways, indicating systematic bias or alternative reasoning mechanism that leads to correct outcomes through unexpected intermediate steps.
@@
  • ($C_{10}$) Reliable oracle. If $\widehat{o_{3}}=o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}=o_{2}$. Observation: Correct prediction with faithful and correct explanation. Interpretation: Strong evidence for expected canonical reasoning pathway, representing the most interpretable and reliable case for knowledge extraction and model understanding.
@@
While this taxonomy relies on the interpreter ability to decode the model’s internal activity, the systematic patterns observed across different models and tasks provide convergent evidence for these behavioral categories. We give examples of thix taxonomy in Appendix H
@@
Table 9,10 and 11 respectively show the detailed impact of hallucination inhibition, linear faithfulness amplification and deceptiveness inhibition steering on self-NLE faithfulness. Categories C2 and C7 are the most prone to be turned into faithful self-NLE overall. Hallucination and faithfulness steering lead to significantly better results as compared to deceptiveness steering overall. Linear faithfulness amplification gives slightly better results than hallucination inhibition. Hallucination inhibition obtains slightly better results than linear faithfulness amplification for C1 and C6.
@@
As shown in Figures 27, 28, and 29, we observe consistent transition patterns when steering makes unfaithful explanations faithful. For initially incorrect predictions, category C1 (complete reasoning failure) consistently transitions to C4 (first-hop reasoning failure), achieving faithfulness approximately 10% of the time under NeuroFaith linear faithfulness amplification and hallucination inhibition. Category C2 (internal-external reasoning mismatch) predominantly transitions to C5 (second-hop reasoning failure) when steering succeeds. Category C3 (explanation-prediction association) exclusively leads to C5 (second-hop reasoning failure) upon becoming faithful.
Initially accurate predictions follow similar patterns. Category C6 (shortcut learning) transitions to C9 (alternative reasoning pathway) when made faithful, while category C8 (explainer parrot) becomes C10 (reliable oracle). Category C7 (deceptiveness or hallucination) can transition to either C9 or C10, though it more commonly becomes a reliable oracle (C10). We give several examples of unfaithful self-NLE made faithful in Appendi H.

### v4

@@ 7 Conclusion
This work introduces NeuroFaith, a framework that measures self-NLE faithfulness by measuring alignment between explanations and mechanistic analysis of model internals. NeuroFaith aligns more closely than existing approaches with the common acceptance of faithfulness. Our findings suggest that faithfulness exhibits linear structure in LLM representation space for 2-hop reasoning and classification tasks, with emergent similarity with other established safety behaviors. Faithfulness linear structure enables both fast detection and enhancement through steering interventions, opening new avenues of research on faithfulness. In this paper, NeuroFaith has been applied on circuit either previously defined or along the last token, based on probe accuracy. It would be valuable to run NeuroFaith based on circuit discovery methods. Our analysis focuses on predict-then-explain scenarios, extending to explain-then-predict settings could reveal how explanation faithfulness relates to performance gains (Bhan et al., 2024). Extending NeuroFaith to chain-of-thought reasoning could also provide valuable comparisons with existing CoT faithfulness studies (Lanham et al., 2023; Turpin et al., 2023).
Impact Statements
This work aims to advance the trustworthiness of AI systems by developing methods to evaluate whether LLM self-explanations actually reflect their internal reasoning processes. Faithful self-explanations are critical for deploying LLMs in high-stakes domains such as healthcare or legal decision-making, where users must understand why a model reached a specific outcome. NeuroFaith provides a principled tool for auditing explanation quality, potentially reducing over-reliance on plausible but misleading justifications. However any potential NeuroFaith user must consider faithfulness score with caution, avoiding over-reliance on a single metric to take high-stake decisions. We especially higlight the importance of human validation during critical NeuroFaith steps such as concept extraction.
Appendix
@@
  • A.
    Scientific Libraries .....................................................................................................................................................................................A
  • B.
    LLM Implementation Details .....................................................................................................................................................................................B
  • C.
    Datasets .....................................................................................................................................................................................C
  • D.
    NeuroFaith Implementation Details .....................................................................................................................................................................................D
    • D.1
      Concept Extraction .....................................................................................................................................................................................D.1
    • D.2
      Mechanistic Interpretation .....................................................................................................................................................................................D.2
    • D.3
      Linear Latent Faithfulness Detection .....................................................................................................................................................................................D.3
  • E.
    Detailed Taxonomy of Self-NLE in Two-hop Reasoning .....................................................................................................................................................................................E
    • E.1
      2-hop reasoning detailed results. .....................................................................................................................................................................................E.1
  • F.
    NeuroFaith Faithfulness Measure Comparison .....................................................................................................................................................................................F
    • F.1
      2-hop Reasoning Faithfulness Qualitative Comparison .....................................................................................................................................................................................F.2
    • F.2
      2-hop Reasoning Faithfulness Quantitative Comparison .....................................................................................................................................................................................F.2
  • G.
    Classification Examples .....................................................................................................................................................................................G
  • H.
    2-hop Reasoning Taxonomy Examples .....................................................................................................................................................................................H
@@ A Scientific Libraries
We used several open-source libraries in this work: pytorch (Paszke et al., 2019), HuggingFace transformers (Wolf et al., 2020) and sklearn (Pedregosa et al., 2011).
@@ C Datasets
Here we provide detailed information about the analyzed datasets. For 2-hop reasoning, we run NeuroFaith on the Wikidata-2-hop dataset (Biran et al., 2024) and Socrates (Yang et al., 2025). We first compute task accuracy on the whole dataset, and then sample 1500 accurate and inaccurate prediction for Wikidata-2-hop. To sample these 1500 instances, we filter out 2-hop reasoning questions where relations are subjective (e.g. ”the most notable work of”) or equivocal, potentially leading to numerous possible answers (e.g. ”the work that features” ). We also sample by setting a maximum number of occurrences for generated answers (15), to foster diversity in the input questions. This filter enables to avoid having too many questions where the answer is a country (e.g. USA). We end up with a dataset made of 3000 samples with 1500 accurate and 1500 inaccurate predictions.
For classification, we run NeuroFaith on AGNews (Gulli, 2005), a newspaper article classification and Ledgar (Tuggener et al., 2020), a more challenging critical domain legal document classification dataset. We retrieve the enriched versions from (Bhan et al., 2025) with labeled concepts for CAV computation. For AGNews, the classes to predict are ’world’, ”sport”, ”business” and ”science & technology”. For Ledgar, the classes to predict are ”Amendments”, ”Survival”, ”Terminations” and ”Terms”. Each dataset is made of 4000 samples.
@@ D.1 Concept Extraction
Here we provide the prompts used to give instructions to Qwen-3-32b to extract relevant concepts (NeuroFaith step 1). For 2-hop reasoning, concept extraction consists in retrieving the bridge object from the self-NLE. Since the input text as the following structures : $x=(o_{1},r_{1},\tiny{\blacktriangle{}},r_{2},\bullet)$, we directly prompt the model to resolve $(o_{1},r_{1},\tiny{\blacktriangle{}}$) by grounding its response on the self-NLE only. We structure our prompt following an in-context learning template with two examples differing from the dataset of interest:
Concept Extraction Prompt for 2-hop Reasoning user
@@
assistant Emmanuel Macron user preprompt + preprompt_example_2 assistant Ingmar Bergman user preprompt + $(o_{1},r_{1},\tiny{\blacktriangle{}}$)
@@
For classification, we assess having access to a set of relevant concepts $\mathcal{C}$ and labels related to the task of interest. Given a certain concept $c$, we only prompt the model to assess if the concept is present in the self-NLE if the concept was initially present in the input text, making the concept extraction process computationally less expensive. We structure our prompt following an in-context learning template with three examples differing from the dataset of interest:
Concept Extraction Prompt for Classification user
@@ D.2 Mechanistic Interpretation
To generate an mechanistic interpretation of an LLM hidden state, the first step is to define the specific neural architectural component and granularity $\mathcal{G}$. Given an index $k$ and a layer $\ell$, a granularity $\mathcal{G}$ refers to a specific point within the model’s computational subgraph where neural activations can be analyzed and from which a local neural interpretation $i^{\ell}_{k}(x)$ is going to be generated. Transformer-based generative LLM can be viewed as a stack of decoder computational blocks (Elhage et al., 2022). The information flow through a single layer and from a layer to another can be described with the following equations:
@@
For the 2-hop reasoning instantiation, the interpreter used to decode $f$ hidden states $\left\{h_{k}^{\ell}\right\}$ is Patchscopes (Ghandeharioun et al., 2024). We use the prompt ”What is the following? Answer briefly [X,X]” to generate the interpretation where X is a token placeholder to be replaced in the latent space by the hidden state to be interpreted. We replace the placeholder tokens at layers 3 and 4 to get two interpretations per hidden state to decode. The layers to be interpreted by Patchscopes vary depending on the assessed model. We set the index token $k$ as the last one related to $o_{1}$ and focus on the late early layers as in (Biran et al., 2024). This way, $\ell\in\left\{{5,6,7,8,9,10,11}\right\}$ for gemma-2-2B, $\ell\in\left\{{8,9,10,11,12,13,14}\right\}$ for gemma-2-9B and $\ell\in\left\{{11,12,13,14,15,16,17}\right\}$ for gemma-2-27B.
@@
Table 5: AGNews concept max. F1 scores ($60\%$) with related layer for gemma-2-27b.
@@
The concept importance $\texttt{I}_{\Gamma}(c)$ can be either computed through representation engineering and concept erasure or gradient-based approaches such as TCAV. We perform concept erasure for $\lambda\in[0,1]$ with a step size of 0.1. Table 5 and 6 show the F1 scores of the properly detected concepts in gemma-2-27b representation space. We estimate the attribution of $c$ given circuit $\Gamma$ as the $\beta_{1}$ parameter of the following regression: $p_{\hat{y}}(x,\left\{\text{do}(H^{\ell}_{k}=h^{\ell}_{k}-\lambda\times\overrightarrow{c_{\ell}})\right\}_{(k,\ell)\in\Gamma})=\beta_{0}+\beta_{1}\times\lambda$. The t-test for $\beta_{1}$ significancy is expressed as follows: $t=\frac{\hat{\beta}_{1}}{SE(\hat{\beta}_{1})}=\frac{\hat{\beta}_{1}}{\sqrt{\frac{MSE}{\sum_{i=1}^{n}(\lambda_{i}-\bar{\lambda})^{2}}}}$, where $\lambda_{i}$ is the $i$ realization of $\lambda$ and $\bar{\lambda}$ is the average $\lambda$ on the analyzed sample.
@@
Faithfulness is finally calculated based on concept importance as follows: $F(x,e)=\frac{1}{p}\sum_{i=1}^{p}\mathbf{1}_{\texttt{I}_{\Gamma}(c_{i})>0}$. Figures 5 and 6 show the faithfulness distributions of NeuroFaith for AGNews and Ledgar for gemma-2-27b.
In Figure 7-12 we plot the concepts sorted by frequency for faithful and unfaithful self-NLE for AGNews and Ledgar for gemma-2-2b and gemma-2-9b for several class predictions. These figures reveal which concepts are associated with faithful versus unfaithful self-NLEs. For instance, Figure 10 shows that NeuroFaith correctly identifies counterintuitive associations (e.g., the concept ”companies” linked to the class ”world”) as unfaithful, demonstrating its reliability.
@@
We also plot the density of faithfulness for AGNews and Ledgar for gemma-2-27b in Figure 5 and 6.
2-hop Reasoning Additional Results.
We show in Table 7 the results obtained by applying NeuroFaith in the case of 2-hop reasoning with Phi-4 as bridge object extractor. The results are highly similar to the ones obtained by using Qwen-3-32B as bridge object extractor (see column ”Faithfulness Corr w/ qwen”).
Model Judge Self-NLE Latent Hop 1 Self-NLE Faithfulness
Correctness Correctness Faithfulness Corr w/ qwen
Accurate Inaccurate Accurate Inaccurate Accurate Inaccurate
gemma-2-2b Phi-4 57.80% 56.50% 47.50% 48.30% 48.00% 56.50% 86.50%
gemma-2-9b Phi-4 73.70% 55.10% 58.50% 44.30% 61.70% 54.80% 91.20%
gemma-2-27b Phi-4 77.80% 55.70% 64.90% 46.90% 69.10% 59.80% 90.90%
Table 7: Self-NLE correctness, first hop latent reasoning correctness, and self-NLE faithfulness across models, evaluated using Phi-4 as judge. ”(In)accurate” represents the set of predictions initially (in)correct. ”Faithfulness Corr w/ qwen” shows faithfulness correlation between faithfulness scores obtained based on Phi-4 and Qwen-3-32B.
Model Task Acc. Self-NLE Latent Hop 1 Self-NLE
Correctness Correctness Faithfulness
Accurate Inaccurate Accurate Inaccurate Accurate Inaccurate
gemma-2-27B 3.4% 72.8% 62.1% 36.4% 8.1% 27.3% 7.6%
Table 8: Self-NLE correctness, first hop latent reasoning correctness, and self-NLE faithfulness for gemma-2-27B on the Socrates dataset.
Figure 13: NeuroFaith faithfulness sensitivity analysis with regards to circuit size, gemma-2-2B.
Figure 14: NeuroFaith faithfulness sensitivity analysis with regards to circuit size, gemma-2-9B.
Figure 15: NeuroFaith faithfulness sensitivity analysis with regards to circuit size, gemma-2-27B.
@@
Here we provide additional information about layer-wise linear probe performance and similarity between faithfulness vectors and with other AI safety linear vectors. Figures 16-24 show the layer-wise performance of the faithfulness linear probes for 2-hop reasoning, AGNews and Ledgar. As shown in Figure 25, 26 and 27, cosine similarity between task-specific linear faithfulness and AI safety behaviors vectors becomes more pronounced with the size of the model.
@@ E Detailed Taxonomy of Self-NLE in Two-hop Reasoning
  • ($C_{1}$) Complete reasoning failure. $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\forall(k,\ell)\in\Gamma,o_{2}\notin\tilde{h^{\ell}_{k}}$. Observation: Wrong prediction, incorrect unfaithful explanation, ground-truth bridge object undetected in circuit $\Gamma$. Interpretation: Evidence suggests failure in first-hop reasoning, with neither explanation nor internal representations containing the expected bridge object.
@@
  • ($C_{2}$) Internal-external reasoning mismatch. $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\exists(k,\ell)\in\Gamma,o_{2}\in\tilde{h^{\ell}_{k}}$. Observation: Wrong prediction, incorrect unfaithful explanation, but ground-truth bridge object detected in circuit $\Gamma$. Interpretation: The model appears to have correct internal knowledge but generates inconsistent explanations, indicating potential reasoning-explanation dissociation because of either deceptiveness or hallucination.
@@
  • ($C_{3}$) Explanation-prediction association. $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=0$ and $\widehat{o_{2}}=o_{2}$. Observation: Wrong prediction with unfaithful but correct explanation. Interpretation: The model associates the correct bridge object with incorrect prediction during explanation generation, suggesting superficial pattern matching without genuinely resolving the first hop of the 2-hop reasoning.
@@
  • ($C_{4}$) First-hop reasoning failure. $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}\neq o_{2}$. Observation: Wrong prediction with faithful but incorrect explanation. Interpretation: The model consistently follows an incorrect reasoning pathway, indicating error in first-hop reasoning or concept misunderstanding.
@@
  • ($C_{5}$) Second-hop reasoning failure. $\widehat{o_{3}}\neq o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}=o_{2}$. Observation: Wrong prediction with faithful and correct explanation. Interpretation: The model correctly identifies he bridge object but fails in second reasoning step, suggesting error occurs after successful first-hop completion.
@@
  • ($C_{6}$) Shortcut learning. $\widehat{o_{3}}=o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\forall(k,\ell)\in\Gamma,o_{2}\notin\tilde{h^{\ell}_{k}}$. Observation: Correct prediction with unfaithful incorrect explanation, ground-truth bridge object undetected. Interpretation: Evidence suggests direct $o_{1}\rightarrow o_{3}$ association, consistent with shortcut learning behavior that bypasses intermediate reasoning steps.
@@
  • ($C_{7}$) Deceptiveness or hallucination. $\widehat{o_{3}}=o_{3}$, $F(x,e)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\exists(k,\ell)\in\Gamma,o_{2}\in\tilde{h^{\ell}_{k}}$. Observation: Correct prediction with unfaithful incorrect explanation, but ground-truth bridge object detected internally. Interpretation: The model possesses correct internal knowledge but generates deceptive (or hallucinated) explanations, suggesting reasoning-explanation dissociation or alternative reasoning pathways. This case is expected to be rare, otherwise highlighting a case where $f$ is not honest in its self-NLE while ”knowing” the ground truth bridge object, raising a problem in $f$ alignment.
@@
  • ($C_{8}$) Explainer parrot. $\widehat{o_{3}}=o_{3}$, $F(x,e)=0$ and $\widehat{o_{2}}=o_{2}$. Observation: Correct prediction with unfaithful but correct explanation. Interpretation: The model generates the expected explanation without corresponding detectable internal reasoning, suggesting post-hoc explanation generation (i.e. ”explainer parrot” behavior).
@@
  • ($C_{9}$) Alternative reasoning pathway. $\widehat{o_{3}}=o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}\neq o_{2}$. Observation: Correct prediction with faithful but incorrect explanation. Interpretation: The model uses a consistent but non-canonical reasoning pathways, indicating bias or alternative reasoning mechanism that leads to correct outcomes through unexpected intermediate steps.
@@
  • ($C_{10}$) Reliable oracle. $\widehat{o_{3}}=o_{3}$, $F(x,e)=1$ and $\widehat{o_{2}}=o_{2}$. Observation: Correct prediction with faithful and correct explanation. Interpretation: Strong evidence for expected canonical reasoning pathway, representing the most interpretable and reliable case for knowledge extraction and model understanding.
@@
While this taxonomy relies on the interpreter ability to decode the model’s internal activity, the systematic patterns observed across different models and tasks provide convergent evidence for these behavioral categories. We give examples of this taxonomy in Appendix H
@@
Table 12,13 and 14 respectively show the detailed impact of hallucination inhibition, linear faithfulness amplification and deceptiveness inhibition steering on self-NLE faithfulness. Categories C2 and C7 are the most prone to be turned into faithful self-NLE overall. Hallucination and faithfulness steering lead to significantly better results as compared to deceptiveness steering overall. Linear faithfulness amplification gives slightly better results than hallucination inhibition. Hallucination inhibition obtains slightly better results than linear faithfulness amplification for C1 and C6.
@@
As shown in Figures 29, 31, and 33, we observe consistent transition patterns when steering makes unfaithful explanations faithful. For incorrect predictions, category C1 (complete reasoning failure) consistently transitions to C4 (first-hop reasoning failure), achieving faithfulness approximately 10% of the time under NeuroFaith linear faithfulness amplification and hallucination inhibition. Category C2 (internal-external reasoning mismatch) predominantly transitions to C5 (second-hop reasoning failure) when steering succeeds. Category C3 (explanation-prediction association) exclusively leads to C5 (second-hop reasoning failure) upon becoming faithful.
Accurate predictions follow similar patterns. Category C6 (shortcut learning) transitions to C9 (alternative reasoning pathway) when made faithful, while category C8 (explainer parrot) becomes C10 (reliable oracle). Category C7 (deceptiveness or hallucination) can transition to either C9 or C10, though it more commonly becomes a reliable oracle (C10). We give several examples of unfaithful self-NLE made faithful in Appendix H.